# PriOr-Flow: Enhancing Primitive Panoramic Optical Flow with Orthogonal View

### v1

@@ 1 Introduction
In the realm of perspective imagery, recent approaches [38, 19, 52, 36, 40, 18] have achieved remarkable success on well-established benchmarks. However, the increasing availability of panoramic cameras has created a pressing demand for panoramic optical flow estimation, which enables comprehensive motion understanding across wide fields of view. Among various projection formats, equirectangular projection (ERP) remains the most prevalent. Yet, by mapping a spherical image onto a 2D plane, ERP introduces severe distortions, particularly near the poles. As a result, applying conventional optical flow methods designed for perspective images directly to panoramic images leads to substantial performance degradation.
@@
To address the limitations of existing ERP-based methods, we introduce PriOr-Flow, a novel dual-branch panoramic optical flow estimation framework that uniquely exploits the low-distortion prior of the orthogonal view to explicitly compensate for severe distortions in polar regions [44, 53, 28, 8]. As noted in [43, 37], the non-uniform sampling of ERP projection results in latitude-dependent distortions following a cosine pattern, with the most severe distortions occurring at the poles (Fig. 3(a)). This distortion significantly degrades optical flow accuracy in these regions. To mitigate this issue, we leverage the orthogonal view, obtained by rotating the panoramic image by $90^{\circ}$ in spherical space (Fig.2). Unlike the primitive ERP view, the orthogonal view exhibits an opposite distortion pattern, featuring minimal distortions in the polar regions (Fig.3). During iterative refinement, PriOr-Flow’s primitive branch adaptively incorporates the low-distortion prior from the orthogonal branch, significantly enhancing optical flow estimation in polar regions.
@@ 2.2 Panoramic Optical Flow Estimation
2.3 Multi-view Panoramic Approaches.
@@ 4.2 Implementation Details
View w/o conf w/ conf EPE SEPE EPE SEPE EPE SEPE
@@
Table 1: Ablation study of the effectiveness of proposed modules. DCCL stands for Dual-Cost Collaborative Lookup, and ODDC represents Ortho-Driven Distortion Compensation. “w/o conf” indicates that we remove the confidence information from the ODDC module. We evaluate the performance of different models in both the equatorial and polar regions.
@@
4.3 Ablation Study
Effectiveness of proposed modules. We progressively incorporate our proposed modules into the baseline model and assess performance across different regions. As shown in Tab. 1, compared to the original RAFT operation, our DCCL module effectively suppresses distortion noise in the cost volume, yielding notable improvements. Adding ODDC—even without incorporating confidence information—still leads to significant gains, primarily benefiting polar regions (an improvement of 25. 3% in the EPE metric). This suggests that motion cues from the orthogonal branch can compensate for distortions in the primitive view. When ODDC is properly integrated with confidence guidance, the complete model (PriOr-RAFT) achieves the best performance, improving accuracy by 26.5%.

### v3

@@ 1 Introduction
In the realm of perspective imagery, recent approaches [41, 22, 55, 39, 8, 43, 21, 9] have achieved remarkable success on well-established benchmarks. However, the increasing availability of panoramic cameras has created a pressing demand for panoramic optical flow estimation, which enables comprehensive motion understanding across wide fields of view. Among various projection formats, equirectangular projection (ERP) remains the most prevalent. Yet, by mapping a spherical image onto a 2D plane, ERP introduces severe distortions, particularly near the poles. As a result, applying conventional optical flow methods designed for perspective images directly to panoramic images leads to substantial performance degradation.
@@
To address the limitations of existing ERP-based methods, we introduce PriOr-Flow, a novel dual-branch panoramic optical flow estimation framework that uniquely exploits the low-distortion prior of the orthogonal view to explicitly compensate for severe distortions in polar regions [47, 56, 31, 11, 10]. As noted in [46, 40], the non-uniform sampling of ERP projection results in latitude-dependent distortions following a cosine pattern, with the most severe distortions occurring at the poles (Fig. 3(a)). This distortion significantly degrades optical flow accuracy in these regions. To mitigate this issue, we leverage the orthogonal view, obtained by rotating the panoramic image by $90^{\circ}$ in spherical space (Fig.2). Unlike the primitive ERP view, the orthogonal view exhibits an opposite distortion pattern, featuring minimal distortions in the polar regions (Fig.3). During iterative refinement, PriOr-Flow’s primitive branch adaptively incorporates the low-distortion prior from the orthogonal branch, significantly enhancing optical flow estimation in polar regions.
@@ 2.2 Panoramic Optical Flow Estimation
2.3 Multi-view Panoramic Approaches
@@ 4.2 Implementation Details
View EPE SEPE EPE SEPE EPE SEPE
@@
Table 1: Ablation study of the effectiveness of proposed modules. DCCL stands for Dual-Cost Collaborative Lookup, and ODDC represents Ortho-Driven Distortion Compensation. We evaluate the performance of different models in both the equatorial and polar regions. Polar regions refer to areas with latitudes greater than $45^{\circ}$, while equatorial regions cover the rest of the sphere.
@@
Model Confidence Poles All
Type EPE SEPE EPE SEPE
Ours w/o conf - 5.65 6.60 3.36 6.37
Ours dist conf Distortion Map 5.62 6.55 3.34 6.34
Ours (PriOr-RAFT) Warp&GW 5.57 6.47 3.30 6.23
Table 2: Ablation study on the type of confidence in ODDC.
4.3 Ablation Study
Effectiveness of proposed modules. We progressively incorporate our proposed modules into the baseline model and assess performance across different regions. As shown in Tab. 1, compared to the original RAFT operation, our DCCL module effectively suppresses distortion noise in the cost volume, yielding notable improvements. Adding ODDC leads to significant gains, primarily benefiting polar regions (an improvement of 26.3% in the EPE metric). This suggests that motion cues from the orthogonal branch can compensate for distortions in the primitive view. We also perform an ablation study on the confidence guidance in ODDC, as shown in Tab. 2. While the distortion map provides view-specific confidence, it is fixed and independent of the input images and optical flow, limiting its ability to reflect actual flow confidence. In contrast, our method explicitly measures feature similarity before and after warping, yielding more accurate confidence estimation and enabling more effective flow fusion.
@@ 5 Conclusion
Acknowledgement. This research is supported by the National Key R&D Program of China (2024YFE0217700), National Natural Science Foundation of China (62472184) and the Fundamental Research Funds for the Central Universities.